# Real-Time Execution of Action Chunking Flow Policies

### v1

@@ 1 Introduction
2 Related Work
Action chunking and VLAs. Inspired in part by human motor control [32], action chunking has recently emerged as the de facto standard in imitation learning for visuomotor control [67, 11]. Learning to generate action chunks from human data requires expressive statistical models, such as variational inference [67, 18], diffusion [11, 12, 68, 67, 45, 58], flow matching [5, 6], vector quantization [33, 3, 43], or byte-pair encoding [46]. Recently, some of these methods have been scaled to billions of parameters, giving rise to VLAs [7, 13, 29, 5, 70, 10, 9, 69, 23, 46, 36], a class of large models built on pre-trained vision-language model backbones. With the capacity to fit ever-growing robot datasets [13, 28, 61, 14, 40, 26], as well as Internet knowledge from vision-language pre-training, VLAs have achieved impressive results in generalizable robot manipulation.
Inpainting and guidance. There is a rich literature on image inpainting with pre-trained diffusion and flow models [47, 54, 39, 41]. In our work, we incorporate one such method [47] into our novel real-time execution framework with modifications (namely, soft masking and guidance weight clipping) that we find necessary for our setting. For sequential decision-making, Diffuser [25] pioneered diffusion-based inpainting for following state and action constraints in long-term planning, though their inpainting method is not guidance-based. Diffuser and other work [63, 1] have also guided diffusion models with value functions to solve reinforcement learning (RL) problems. Our work is distinct in that it is the first to apply either inpainting or guidance to real-time control.
Real-time execution. Real-time control has been studied long before the advent of VLAs. Similar to action chunking, model predictive control (MPC; [50]) generates plans over a receding time horizon; like our method, it parallelizes execution and computation, and uses the prior chunk to warm-start planning for the next. Though recent works combining learning methods with MPC have demonstrated real-time control capabilities in narrow domains [52, 20], their reliance on explicit dynamic models and cost functions makes their application difficult in unstructured settings, where VLAs have gained prominence. Separately, in reinforcement learning, a variety of prior works have developed time-delayed decision-making methods [56, 15, 53, 62, 65, 66]. However, these approaches are not always applicable to imitation learning, and none of them leverage action chunking. Most recently, hierarchical VLA designs [57, 4] have emerged where the model is split into a System 2 (high-level planning) and System 1 (low-level action generation) component. The System 2 component contains the bulk of the VLA’s capacity and runs at a low frequency, while the System 1 component is lightweight and fast. This approach is orthogonal to ours, and comes with its own tradeoffs (e.g., limiting the size of the System 1 component and requiring its own training recipe).
@@ 3 Preliminaries and Motivation
We begin with an action chunking policy denoted by $\pi(\mathbf{A}_{t}|\mathbf{o}_{t})$, where $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H-1}]$ is a chunk of future actions, $\mathbf{o}_{t}$ is an observation, and $t$ indicates a controller timestep. We call $H$ the prediction horizon. When action chunking policies are rolled out, the first $s\leq H$ actions from each chunk are executed, and then a new chunk is produced by the policy. We call $s$ the execution horizon; often it is shorter than the prediction horizon, but still much greater than 1 (e.g., $s\approx H/2$  [11, 5, 23]). Chunked execution ensures temporal consistency at the expense of reactivity. A long execution horizon reduces a policy’s responsiveness to new information, while a short one increases the likelihood of mode-jumping, jerky behavior resulting from discontinuities between chunks.
@@
Now, let $\Delta t$ be sampling period of a controller, i.e., the duration of a controller timestep, and let $\delta$ be the time it takes for the policy to generate an action chunk. We define $d:=\lfloor\delta/\Delta t\rfloor$ and call this quantity the inference delay, corresponding to number of controller timesteps between when $\mathbf{o}_{t}$ is received and when $\mathbf{A}_{t}$ is available. [^1] For simplicity, we do not consider delays or synchronization issues at the sub-timestep level; we assume that the low-level controller provides $\mathbf{o}_{t}$ at the same instant that $\mathbf{a}_{t-1}$ is consumed. If $d=0$, then inference can be performed between two timesteps without any interruption. However, this is near impossible to achieve with modern VLAs. For example, with an NVIDIA RTX 4090 GPU, the 3 billion parameter $\pi_{0}$ VLA spends 46ms on the KV cache prefill alone, before any denoising steps [5], and targets a 50Hz control frequency ($\Delta t=20$ms). Run in remote inference for mobile manipulation, $\pi_{0}$ lists 13ms of network latency, in perfect conditions with a wired connection. In a more realistic setting, the network overhead alone could easily exceed 20ms. Kim et al. [30], who optimize the 7B OpenVLA model [29] specifically for inference speed, achieve no better than 321ms of latency on an NVIDIA A100 GPU.
When $d>0$, chunked execution becomes more complex. Naive synchronous inference, the default in many prior works [5, 29, 8, 23, 30, 58], introduces visible pauses between chunks that not only slow down execution but also change the dynamics of the robot, introducing distribution shift between training and evaluation. The first requirement of a real-time system is asynchronous execution, where inference is started early so that an action is guaranteed to be available at every timestep.
Let $\mathbf{a}_{t^{\prime}|t}$ denote the $(t^{\prime}-t)$-th action of chunk $\mathbf{A}_{t}$, generated from observing $\mathbf{o}_{t}$. If $\mathbf{A}_{0}$ is currently executing, and we wish to switch chunks at step $j$, then an asynchronous algorithm must start inference at $j-d$. However, since the policy cannot know what will happen between steps $j-d$ and $j$ while generating $\mathbf{A}_{j-d}$, the transition point between $\mathbf{a}_{j-1|0}$ and $\mathbf{a}_{j|j-d}$ may be arbitrarily discontinuous and out-of-distribution. Similar to a too-short execution horizon, this strategy leads to jerky behavior that is worsened dramatically with higher delays; see Figure 2.
@@ 4.1 Inference-Time Inpainting with Flow Matching
Inpainting is a known strength of iterative denoising frameworks such as diffusion and flow matching. We build on the training-free image inpainting algorithm from Pokle et al. [47], which is itself based on pseudoinverse guidance ($\Pi$GDM; [54]). The algorithm operates by adding a gradient-based guidance term to the learned velocity field $\mathbf{v}$ at each denoising step (Equation 1) that encourages the final generation to match some target value, $\mathbf{Y}$, which is a corrupted version of the desired result. In the case of image inpainting, the corruption operator is masking, $\mathbf{Y}$ is the masked image, and the desired result is a full image consistent with $\mathbf{Y}$ in the non-masked areas. The $\Pi$GDM gradient correction, specialized to our setting, is given by
@@
$\widehat{\mathbf{A}^{1}_{t}}$ is an estimate of the final, fully denoised action chunk and $\mathbf{W}$ is the mask. We are abusing notation by treating $\mathbf{Y}$, $\mathbf{A}_{t}$, and $\mathbf{W}$ as vectors of dimension $HM$ where $M$ is the dimension of each action. Thus, the guidance term is a vector-Jacobian product and can be computed using backpropagation. The guidance weight clipping, $\beta$, is our addition; we found that without it, the algorithm became unstable with the small number of denoising steps commonly used in control problems (see A.2 for ablation).
@@ 4.2 Soft Masking for Improved Cross-Chunk Continuity
In practice, naively inpainting using only the first $d$ timesteps of the previous action chunk is often insufficient to ensure that the new chunk takes a consistent strategy, particularly when $d$ is small (e.g., see Figure 4). The $\Pi$GDM correction is not perfect, and a small $d$ leads to a weak guidance signal, which can allow for the new chunk to still switch strategies and cause discontinuities. Our solution, illustrated in Figure 3, is to give our policy more cross-chunk continuity by considering not just the first $d$ overlapping actions, but all $H-s$ overlapping actions. We do this via soft masking, setting $\mathbf{W}$ to real-valued weights rather than 1s and 0s. The first $d$ actions get a weight of 1; the last $s$ actions of the new chunk do not overlap with the previous chunk, so they get a weight of 0; the actions in between get weights that exponentially decay from 1 to 0, accounting for the fact that actions further in the future should be treated with more uncertainty. The resulting expression for $\mathbf{W}$ is given by
@@
Intuitively, $\mathbf{W}$ modulates the “attention” paid to each corresponding action from the previous chunk.
@@ 4.3 Real-Time Chunking
12:    $\mathcal{Q}=$ new Queue([$d_{\text{init}}$], maxlen=$b$) $\triangleright$ Holds a limited buffer of past inference delays
@@
15:        $s=t$triangleright $s$ is the number of actions executed since last inference started
16:        $\mathbf{A}_{\text{prev}}=\mathbf{A}_{\text{cur}}[s,s+1,\ldots,H-1]$triangleright Remove the $s$ actions that have already been executed
@@
18:        $d=$ max$(\mathcal{Q})$triangleright Estimate the next inference delay conservatively
@@
20:           $\mathbf{A}_{\text{new}}=$ GuidedInference$(\pi,\mathbf{o},\mathbf{A}_{\text{prev}},d,s)$
21:        $\mathbf{A}_{\text{cur}}=\mathbf{A}_{\text{new}}$triangleright Swap to the new chunk as soon as it is available
22:        $t=t-s$triangleright Reset $t$ so that it indexes into $\mathbf{A}_{\text{new}}$
23:        enqueue $t$ onto $\mathcal{Q}$triangleright Record the observed delay
@@
27:        $f_{\widehat{\mathbf{A}^{1}}}=\mathbf{A}^{\prime}\mapsto\mathbf{A}^{\prime}+(1-\tau)\mathbf{v}_{\pi}(\mathbf{A}^{\prime},\mathbf{o},\tau)$triangleright Define denoising function (Eq. 3)
28:        $\mathbf{e}=\left(\mathbf{A}_{\text{prev}}-f_{\widehat{\mathbf{A}^{1}}}(\mathbf{A}^{\tau})\right)^{\top}\operatorname{diag}(\mathbf{W})$triangleright Weighted error term from Eq. 2
29:        $\mathbf{g}=\mathbf{e}\cdot\left.\frac{\partial f_{\widehat{\mathbf{A}^{1}}}}{\partial\mathbf{A}^{\prime}}\right|_{\mathbf{A}^{\prime}=\mathbf{A}^{\tau}}$triangleright Compute vector-Jacobian product from Eq. 2 via autodiff
30:        $\mathbf{A}^{\tau+\frac{1}{n}}=\mathbf{A}^{\tau}+\frac{1}{n}\left(\mathbf{v}_{\pi}(\mathbf{A}^{\tau},\mathbf{o},\tau)+\min\left(\beta,\frac{1-\tau}{\tau\cdot r^{2}_{\tau}}\right)\mathbf{g}\right)$triangleright Integration step (Eq. 1)     return $\mathbf{A}^{1}$
@@ 5 Experiments
We first evaluate RTC using a benchmark of 12 highly dynamic and stochastic environments in the Kinetix [42] simulator. We use this benchmark to compare the performance of RTC to other methods under simulated inference delays, as well as investigate the effect of soft masking. Then, using the $\pi_{0.5}$ VLA [23] as the base model, we evaluate the performance and speed of RTC on 6 challenging bimanual dexterous manipulation tasks, including 2 mobile manipulation tasks (see Figure 6).
@@ 5.2 Real-World Results
Next, we deploy our full real-time chunking system to the real world. We use the $\pi_{0.5}$ [23] as our base policy, and evaluate RTC on a bimanual system with two 6-DoF arms and parallel jaw grippers. Unlike our simulated benchmark, the robots use position control, and so synchronous inference—stopping between chunks—is a reasonable default strategy, used in many prior works [5, 23, 30, 46]. Our goal is to improve upon synchronous inference in a combination of both performance and speed.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: The six real-world tasks, including 2 mobile manipulation tasks. See the blog post for videos.
@@
See Figure 6 for images of tasks, and the blog post for videos. We evaluate each task and method for 10 trials for a total of 480 episodes, adding up to 28 hours of pure robot execution time. We also post-hoc annotate the score for each episode and the timestamp at which each step is achieved.
@@
  • TE, sparse. This is similar to naive async in our simulated results; it executes $s=25$ actions at a time while computing the next chunk in parallel. We found it significantly reduced jerkiness to also apply TE, even though only the first $H-s-2d$ executed steps of each chunk have overlapping actions to ensemble.
  • TE, dense. This strategy is the closest to the original TE in Zhao et al. [67]. We run inference as often as possible by setting $s_{\text{min}}=1$ in Algorithm 1, meaning $s=d$ for every chunk. This results in there always being at least 2 overlapping action chunks to ensemble, and often more.
@@
Figure 7: Top: Controller steps (equivalent to elapsed time with inference pauses removed multiplied by 50Hz) vs. cumulative progress for each task, aggregated across all delays. Progress is measured in discrete steps corresponding to the subsections of each task. Left: Time (including inference pauses) vs. cumulative progress aggregated across all tasks. The x-axis is log scale to better show progress during both short and long-horizon tasks. Right: Inference delay vs. average throughput, defined as the proportion of task completed divided by duration of episode averaged over episodes. Error bars are ±$\pm 1$ SEM. Average throughput gives a balanced view of both speed and performance for each method. Neither TE variant can run at +100 or +200ms of injected latency, causing such high oscillations that the robot’s protective stop is triggered.

### v2

@@ 2 Preliminaries and Motivation
We begin with an action chunking policy denoted by $\pi(\mathbf{A}_{t}|\mathbf{o}_{t})$, where $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H-1}]$ is a chunk of future actions, $\mathbf{o}_{t}$ is an observation, and $t$ indicates a controller timestep. We call $H$ the prediction horizon. When action chunking policies are rolled out, only the first $s\leq H$ actions from each chunk are executed. We call $s$ the execution horizon; often it is shorter than the prediction horizon, but still much greater than 1 (e.g., $s\approx H/2$  [11, 5, 24]). Chunked execution ensures temporal consistency at the expense of reactivity. A long execution horizon reduces a policy’s responsiveness to new information, while a short one increases the likelihood of mode-jumping, jerky behavior resulting from discontinuities between chunks.
@@
Now, let $\Delta t$ be sampling period of the controller, i.e., the duration of a controller timestep, and let $\delta$ be the time it takes for the policy to generate an action chunk. We say that a system is real-time if it is guaranteed to produce a response (in our case: $\mathbf{a}_{t}$) to an event (receiving $\mathbf{o}_{t}$) within a fixed time constraint ($\Delta t$). If $\delta\leq\Delta t$, then meeting the real-time constraint is trivial, since an entire chunk can be generated between two controller timesteps. However, this is near impossible to achieve with modern VLAs. For example, with an RTX 4090 GPU, the 3 billion parameter $\pi_{0}$ VLA spends 46ms on the KV cache prefill alone, before any denoising steps [5], and targets a 50Hz control frequency ($\Delta t=20$ms). Run in remote inference for mobile manipulation, $\pi_{0}$ lists 13ms of network latency, in perfect conditions with a wired connection. In a more realistic setting, the network overhead alone could easily exceed 20ms. Kim et al. [31], who optimize the 7B OpenVLA model [30] specifically for inference speed, achieve no better than 321ms of latency on a server-grade A100 GPU.
Naive synchronous inference, the default in many prior works [5, 30, 8, 24, 31, 59], simply starts inference at the end of the execution horizon and waits while the policy generates the next chunk. When $\delta>\Delta t$, this introduces visible pauses between chunks that not only slow down execution but also change the dynamics of the robot, introducing distribution shift between training and evaluation. To develop a real-time strategy, we must first introduce asynchronous inference, where inference is started early and happens concurrently with execution.
We define $d:=\lfloor\delta/\Delta t\rfloor$ and call this quantity the inference delay, corresponding to number of controller timesteps between when $\mathbf{o}_{t}$ is received and when $\mathbf{A}_{t}$ is available. [^1] For simplicity, we do not consider delays or synchronization issues at the sub-timestep level; we assume that the environment or lower-level controller provides $\mathbf{o}_{t}$ at the same instant that $\mathbf{a}_{t-1}$ is consumed. Let $\mathbf{a}_{t^{\prime}|t}$ denote the $(t^{\prime}-t)$-th action of chunk $\mathbf{A}_{t}$, generated from observing $\mathbf{o}_{t}$. If $\mathbf{A}_{0}$ is currently executing, and we desire an execution horizon of $s$, then an asynchronous algorithm must start inference at sds-d. So long as $d\leq H-s$, then this strategy will satisfy the real-time constraint and guarantee that an action is always available when it is needed. However, since the policy cannot know what will happen between steps sds-d and $s$ while generating $\mathbf{A}_{s-d}$, the transition point between $\mathbf{a}_{s-1|0}$ and $\mathbf{a}_{s|s-d}$ may be arbitrarily discontinuous and out-of-distribution. Similar to a too-short execution horizon, this strategy leads to jerky behavior that is worsened dramatically with higher delays; see Figure 2.
@@ 3.1 Inference-Time Inpainting with Flow Matching
[…]
@@
$\widehat{\mathbf{A}^{1}_{t}}$ is an estimate of the final, fully denoised action chunk and $\mathbf{W}$ is the mask. We are abusing notation by treating $\mathbf{Y}$, $\mathbf{A}_{t}$, and $\mathbf{W}$ as vectors of dimension $HM$ where $M$ is the dimension of each action. Thus, the guidance term is a vector-Jacobian product and can be computed using backpropagation. The guidance weight clipping, $\beta$, is our addition; we found that without it, the algorithm became unstable with the small number of denoising steps commonly used in control problems (see A.2 for an ablation).
@@ 3.2 Soft Masking for Improved Cross-Chunk Continuity
[…]
@@
Intuitively, $\mathbf{W}$ modulates the “attention” paid to each corresponding action from the previous chunk. See Appendix A.4 for a comparison between different decay schedules.
@@ 3.3 Real-Time Chunking
12:  $\mathcal{Q}=$ new Queue($d_{\text{init}}$], maxlen=$b$) $\triangleright$ Holds a limited buffer of past inference delays
@@
15:   $s=t$ $\triangleright$ $s$ is the number of actions executed since last inference started
16:   $\mathbf{A}_{\text{prev}}=\mathbf{A}_{\text{cur}}[s,s+1,\ldots,H-1]$ $\triangleright$ Remove the $s$ actions that have already been executed
@@
[…]
@@
[…]
21:   $\mathbf{A}_{\text{cur}}=\mathbf{A}_{\text{new}}$ $\triangleright$ Swap to the new chunk as soon as it is available
22:   t=tst=t-s $\triangleright$ Reset $t$ so that it indexes into $\mathbf{A}_{\text{new}}$
23:   enqueue $t$ onto $\mathcal{Q}$ $\triangleright$ Record the observed delay
@@
27:   $f_{\widehat{\mathbf{A}^{1}}}=\mathbf{A}^{\prime}\mapsto\mathbf{A}^{\prime}+(1-\tau)\mathbf{v}_{\pi}(\mathbf{A}^{\prime},\mathbf{o},\tau)$ $\triangleright$ Define denoising function (Eq. 3)
28:   $\mathbf{e}=\left(\mathbf{A}_{\text{prev}}-f_{\widehat{\mathbf{A}^{1}}}(\mathbf{A}^{\tau})\right)^{\top}\operatorname{diag}(\mathbf{W})$ $\triangleright$ Weighted error term from Eq. 2
29:   $\mathbf{g}=\mathbf{e}\cdot\left.\frac{\partial f_{\widehat{\mathbf{A}^{1}}}}{\partial\mathbf{A}^{\prime}}\right|_{\mathbf{A}^{\prime}=\mathbf{A}^{\tau}}$ $\triangleright$ Compute vector-Jacobian product from Eq. 2 via autodiff
30:   $\mathbf{A}^{\tau+\frac{1}{n}}=\mathbf{A}^{\tau}+\frac{1}{n}\left(\mathbf{v}_{\pi}(\mathbf{A}^{\tau},\mathbf{o},\tau)+\min\left(\beta,\frac{1-\tau}{\tau\cdot r^{2}_{\tau}}\right)\mathbf{g}\right)$ $\triangleright$ Integration step (Eq. 1)   return $\mathbf{A}^{1}$
@@ 4 Experiments
We first evaluate RTC using a benchmark of 12 highly dynamic and stochastic environments in the Kinetix [43] simulator. We use this benchmark to compare the performance of RTC to other methods under simulated inference delays, as well as investigate the effect of soft masking. Then, using the $\pi_{0.5}$ VLA [24] as the base model, we evaluate the performance and speed of RTC on 6 challenging bimanual dexterous manipulation tasks, including 2 mobile manipulation tasks.
@@ 4.2 Real-World Results
Next, we deploy our full real-time chunking system to the real world. We use the $\pi_{0.5}$VLA [24] as our base policy, and evaluate RTC on a bimanual system with two 6-DoF arms and parallel jaw grippers. Unlike our simulated benchmark, the robots use position control, and so synchronous inference—stopping between chunks—is a reasonable default strategy, used in many prior works [5, 24, 31, 47]. Our goal is to improve upon synchronous inference in a combination of both performance and speed.
@@
See the accompanying blog post for videos of each task. We evaluate each task and method for 10 trials for a total of 480 episodes, adding up to 28 hours of pure robot execution time. We also post-hoc annotate the score for each episode and the timestamp at which each step is achieved.
@@
  • TE, sparse. This is similar to naive async in our simulated results; it executes $s=25$ actions at a time while computing the next chunk in parallel. We found it significantly reduced jerkiness to also apply TE, even though only the first Hs2dH-s-2d executed steps of each chunk have overlapping actions to ensemble.
  • TE, dense. This strategy is the closest to the original TE in Zhao et al. [68]. We run inference as often as possible, resulting in $s=d$ for every chunk. This results in there always being at least 2 overlapping action chunks to ensemble, and often more.
@@
Figure 6: Top: Controller steps (equivalent to elapsed time with inference pauses removed multiplied by 50Hz) vs. cumulative progress for each task, aggregated across all delays. Progress is measured in discrete steps corresponding to the subsections of each task. Left: Time (including inference pauses) vs. cumulative progress aggregated across all tasks. The x-axis is log scale to better show progress during both short and long-horizon tasks. Right: Inference delay vs. average throughput, defined as the proportion of task completed divided by duration of episode averaged over episodes. Error bars are $\pm 1$ SEM. Average throughput gives a balanced view of both speed and performance for each method. Neither TE variant can run at +100 or +200ms of injected latency, causing such high oscillations that the robot’s protective stop is triggered.
@@
5 Related Work
Action chunking, VLAs, and cascade control. Inspired in part by human motor control [33], action chunking has recently emerged as the de facto standard in imitation learning for visuomotor control [68, 11]. Learning to generate action chunks from human data requires expressive statistical models, such as variational inference [68, 19], diffusion [11, 12, 69, 68, 46, 59], flow matching [5, 6], vector quantization [34, 3, 44], or byte-pair encoding [47]. Recently, some of these methods have been scaled to billions of parameters, giving rise to VLAs [7, 13, 30, 5, 71, 10, 9, 70, 24, 47, 37], a class of large models built on pre-trained vision-language model backbones. With the capacity to fit ever-growing robot datasets [13, 29, 62, 15, 41, 27], as well as Internet knowledge from vision-language pre-training, VLAs have achieved impressive results in generalizable robot manipulation. When applied to real-world robots, action chunking policies are often used in conjunction with a lower-level, higher-frequency control loop—such as a PID controller—which translates the outputs of the policy (e.g., joint positions) to hardware-specific control signals (e.g., joint torques). In these cases, action chunking policies can be viewed as a form of cascade control [14], with the learned policy acting the outermost control loop. However, this is not always the case: for example, our simulated experiments use learned policies that output torques and forces directly. As such, we defer any exploration of the intersection between cascade control theory and learned action chunking policies to future work.
Inpainting and guidance. There is a rich literature on image inpainting with pre-trained diffusion and flow models [48, 55, 40, 42]. In our work, we incorporate one such method [48] into our novel real-time execution framework with modifications (namely, soft masking and guidance weight clipping) that we find necessary for our setting. For sequential decision-making, Diffuser [26] pioneered diffusion-based inpainting for following state and action constraints in long-term planning, though their inpainting method is not guidance-based. (See Appendix A.4 for a comparison to the inpainting method from Diffuser applied to our setting.) Diffuser and other work [64, 1] have also guided diffusion models with value functions to solve reinforcement learning (RL) problems. Our work is distinct in that it is the first to apply either inpainting or guidance to real-time control.
Real-time execution. Real-time control has been studied long before the advent of VLAs. Similar to action chunking, model predictive control (MPC; [51]) generates plans over a receding time horizon; like our method, it parallelizes execution and computation, and uses the prior chunk to warm-start planning for the next. Though recent works combining learning methods with MPC have demonstrated real-time control capabilities in narrow domains [53, 21], they rely on explicit, hand-crafted dynamics models and cost functions. These methods are not applicable to our setting, which considers model-free imitation learning policies and tests them on unstructured, open-world manipulation tasks. Separately, in reinforcement learning, a variety of prior works have developed time-delayed decision-making methods [57, 16, 54, 63, 66, 67]. However, these approaches are not always applicable to imitation learning, and none of them leverage action chunking. Most recently, hierarchical VLA designs [58, 4] have emerged where the model is split into a System 2 (high-level planning) and System 1 (low-level action generation) component. The System 2 component contains the bulk of the VLA’s capacity and runs at a low frequency, while the System 1 component is lightweight and fast. This approach is orthogonal to ours, and comes with its own tradeoffs (e.g., limiting the size of the System 1 component and requiring its own training recipe).